Q. A diffraction grating has 3 x 106 lines per meter. The grating is illuminated by monochromatic plane waves of wavelength 600 nm at normal incidence that forms an interference pattern on a screen located 0.50 m away

a) Find the angle the first bright fringe.

b) Find the angle corresponding to the second bright fringe.

c) Find the angle of the third bright fringe.
Answer #1

Slit width, d 3.33 * 10-7m n 3 * 106

Location of Maxima ( bright fringe ) is : -1 0 = sin mX d

a. For first bright fringe, m = 0

B = sin -1 0 * 600 * 10 3.33* 10-7 = 0° → Answer

b. For second bright fringe, m = 2

-1 >> = sin 1 * 600 * 10 3.33 * 10-7 = 10.4° → Answer

c. For third bright fringe, m =3

-1 >02 = sin 2 * 600 * 10 3.33* 10-7 = 21.1° → Answer